OpenStudy (anonymous):

For a triangle with sides a, b and c, what is the range of the length of side c? a = 418, b = 222, c =

OpenStudy (anonymous):

Pythagorean theorem

OpenStudy (anonymous):

which is...

OpenStudy (anonymous):

a ^2 + b^2 = c^2

OpenStudy (anonymous):

so you would do 418^2 + 222^2 = c^2 then take those values and square them then add them and then take the square root of both sides that should tell you what C is

OpenStudy (anonymous):

theres two answers

OpenStudy (anonymous):

there's really only one were you given a different formula in class, b/c if you're not advanced, then you probably havent used pythagorean

OpenStudy (anonymous):

actually, they probably want you to only use whole numbers: therefore the range would be between 473 and 474

OpenStudy (anonymous):

the question is For a triangle with sides a, b and c, what is the range of the length of side c? a = 418, b = 222, c = __ < c <__

OpenStudy (anonymous):

http://www.algebra.com/algebra/homework/Triangles/Triangles.faq.question.253712.html this site might explain it better

OpenStudy (anonymous):

thankd

OpenStudy (anonymous):

thanks*

OpenStudy (anonymous):

so does anybody know what the answer is to this question?